This morning, my dog Walter broke out of the house and ran directly to the farm up the road at a speed of 12mph. When he arrived at the farm he was scared by the cows so he turned around and ran directly home (by the same route) at a speed of 8mph. Walter spent a total time of one hour running to the farm and back. Find the distance from my house to the farm.
Answer by nerdybill(5411)

.
Let t = time to get to the farm
then
1-t = time to get back to the house
.
Applying the distance formula:
d = rt
where
d is distance
r is rate or speed
t is time
.
12t = 8(1-t)
12t = 8-8t
20t = 8
t = 8/20
t = 2/5 hour (to run from house to farm)
.
Distance then is:
d = (2/5)12
d = 24/5 = 4.8 miles (distance from house to farm)